A restaurant bought on the first of January 400 bottles of vine at a price of 40 DKK per bottle.
Payment condition was current month + 15 days. After the first quarter it appeared that the sales in January were 100 bottles, sales in February 125 bottles and the sales in March were 175 bottles .All bottles of wine were sold at a price of 150 DKK each.
Prepare a profit and loss statement
1
Expert's answer
2020-04-05T12:46:03-0400

January February March 1st quarter

Costs 400*40=16 000 0 0 16 000

Revenue 100*150=15 000 125*150=18 750 175*150=26 250 60 000

profit , loss 15 000-16000=-1000 18 750-0=18 750 26 250-0=26 250 60 000-16 000=44 000


based on the results of the 1st quarter, profit in the amount of 44 000 DKK